Lois rode her bike to visit a friend. She traveled 10 mi/h. While she was there, it started to rain. Her friend dorve her home in a car traveling 20 mi/h. Lois took 1.5 h longer to go to her friend's than to return home. How many hours did it take Lois to ride to her friends house?

MY WORK-
r t d
10 mi/h 1.5+t 10(1.5+t)
20 mi/h t 20t
10(1.5+t)=20t
10(1.5+t)=20t
15+10t=20t
15+10t-10t=20t
15=10t

Answer by mananth(16946)

Lois rode her bike to visit a friend. She traveled 10 mi/h. While she was there, it started to rain. Her friend dorve her home in a car traveling 20 mi/h. Lois took 1.5 h longer to go to her friend's than to return home. How many hours did it take Lois to ride to her friends house?
...
let the time to friends house be x hours
speed = 10 mph
distance = 10x
...
return
speed = 2o mph
time = x-1.5
distance = 20*(x-1.5)
...
10x=20(x-1.5)
10x=20x-30
10x-20x=-30
-10x=-30
/-10
x= 3 hours
...
m.ananth@hotmail.ca